OpenStudy (danjs):

Solve: y'' + y = csc(x) ; 0

OpenStudy (danjs):

general soln

OpenStudy (loser66):

characteristic equation is??

OpenStudy (danjs):

variation of parameters

OpenStudy (loser66):

nope, solve it by characteristic equation.

OpenStudy (danjs):

For the homogeneous solution to L(y) = 0 characteristic eq is.. r^2 + 1 = 0 r = + or - i \[y _{h} = c _{1}e^{i * x} + c _{2}e^{-i * x}\]

OpenStudy (danjs):

been a few years, bear with me, looking at the book. lol

OpenStudy (loser66):

is it not that if \(r=\pm i\) , then \(y_h= C_1 cos (x) + C_2 sin(x)\)??

OpenStudy (danjs):

so you can say...from Euler

OpenStudy (danjs):

right

OpenStudy (loser66):

ok, now solve for partial part.

OpenStudy (danjs):

power series expansion of e^(ix)

OpenStudy (danjs):

Particular solution to L(y) = f(x), ... hmm

OpenStudy (danjs):

minute

OpenStudy (danjs):

looking at variation of parameter technique

OpenStudy (danjs):

\[y _{p} = g _{1}u _{1} + g _{2}u _{2}\] so, u1 = cos(x) u2 = sin(x)

OpenStudy (danjs):

and \[g _{1}^{~'}*u _{1} + g _{2}^{~'}*u _{2} = 0\] \[g_{1}^{'}*u _{1}^{'} + g_{2}^{'}*u _{2}^{'} = f(x)\] so just put in u1 and u2, and solve g1 and g2 prime ?

OpenStudy (danjs):

then integrate to get the g1 and g2 for the particular solution...?

OpenStudy (danjs):

i remember memorizing those 2 lines , forgot where they come from, the = 0 and = f(x) , things

OpenStudy (loser66):

I am sorry, I am not familiar with this method. I use Wronskian to solve it. :) @ganeshie8 Please.

OpenStudy (danjs):

ah right, i did not do that one

OpenStudy (danjs):

so i got \[g _{2} = \ln(\sin(x)) ~~~so~~~~g _{1} = -x\] then, \[y _{p} = -x*\cos(x) + \ln(\sin x)*\sin(x)\]

OpenStudy (danjs):

u get that with your method?

OpenStudy (danjs):

it has to be right, it cancelled nicely, and is a book 'nice' problem.

OpenStudy (danjs):

\[y _{g} = y _{h} + y _{p}\]

ganeshie8 (ganeshie8):

looks good, wolfram agrees :) http://www.wolframalpha.com/input/?i=solve+y%27%27+%2B+y+%3D+csc%28x%29++++++++++

OpenStudy (danjs):

cool, trying to do a couple of each solution technique to review, i forget a F ton